The oxidation of methanol, as described by the equ
Answer #1

2CH3OH+3O2------> 2Co2 + 3H2O

carbon oxidation number increases from -2 to +4 . total 12 electrons are involve

\DeltaG0   = -nE0 *F

-937900 =-12*E0 *96500

-937900 = -E0 *1158000

E0 = 937900/1158000 = 0.81V